Grade 11Trigonometry

Write the area of the triangle formed by the coordinate axis and the line (sec theta-tan theta)x+(sec theta+tan theta)y=2

Sunit Biswal
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
(sec theta – tan theta) x  + (sec theta + tan theta) y = 2
 
(sec theta – tan theta)x /2  + (sec theta + tan theta) y/2 = 1
 
hence area = ½ * 2/ (sec theta – tan theta) * 2/ (sec theta + tan theta)
 
 = 2 / (sec^2 theta – tan^2 theta) = 2